OpenStudy (anonymous):

solve the equation= square root of x-3= x-3

OpenStudy (anonymous):

square both sides and you should get (x-3) = (x-3)^2. (x-3)^2 = (x)^2 -6x+9.

OpenStudy (anonymous):

minus x and -3 from both sides and you should get (x)^2-7x+12. which equals (x-3)(x-4). x=3 and 4